# CoPlay: Audio-agnostic Cognitive Scaling for Acoustic Sensing

### v1

@@ 4.2. Experiment results
Figure 6. Ablation study results.
Generalization across datasets and music types: In total, we have 8 datasets, with two datasets for each of the four audio types: piano, hot songs, bass, and speech. When the model is trained and tested with the same dataset, we refer to it as intra-dataset experiment. When training and testing with different datasets from the same type, it is a cross-dataset experiment. Lastly, to assess a more out-of-distribution case, we train on one dataset from one type and test on another dataset from another type, termed as cross-type experiment. Note that, in intra-dataset experiments, we avoid randomly splitting the dataset for testing that is non-realistic in usage. Instead, we take the last 10% of the video as testing set. The rest is then randomly split into 80% training and 20% validation. Fig. 5 shows the average error for an extensive run of these experiments. The results for both sine and chirps are separately plotted as bar charts for different groups of music types. To better plot the error, the y-axis is log-scaled testing loss multiplied by $51e$ 5.
In conclusion, the model manifests the capability of generalizing to different types of music or speech, since the cross-dataset error is similar to that of cross-type data error for both sine and FMCW chirps. While intra-dataset usually outperforms the others because the distribution between testing and training is close, the intra-dataset results in chirp are a bit counter-intuitive and it takes 10x more epochs to converge. We think the reason could be that chirp is more complex and needs more data to train, while intra-dataset trains on only 6̃0% of cross-dataset/cross-type experiments. Besides, it is worth mentioning that the cross-type results in the speech group are also good. This is a particularly significant generalization result, since the speech category stands out as the most distinguishable type of audio, compared with music of piano, hot songs, and bass.
target -loss recovery -loss amplitude -loss variance -loss
sine 0.5928 0.0012 0.9558 -
chirp 0.1852 0.0013 0.9560 0.0214
Table 1. The value of sub-loss terms.
Decompose the loss: The overall error mentioned above comprises multiple loss terms as defined in the section 3.3. Hence, we decompose it into individual loss term values to assess the output per each objective represented by the corresponding loss term. Table. 1 demonstrates the average sub-loss on hot songs datasets. They are training loss including three sub-loss values for sine and an additional variance-loss for FMCW, which we defined above to even the amplitude in the target bandwidth. By default weights, each sub-loss ranges from 0 to 1. We observe that target-loss and amplitude-loss are the bottlenecks, aligning with the intuition that minimizing frequency distortion and maximizing the amplitude are opposed objectives, which is a common issue in multi-task learning. This table serves as a reference for tweaking the weights of loss terms to specific downstream task requirements. For instance, near-field sensing tasks can sacrifice amplitude for minimal distortion.
Varying music volume: Although it is possible that a high volume of music might degrade our model performance, Fig. 5(b) shows that the overall error does not increase with music volume. In detail, we train with the same sub-loss experiment and test with increasing levels of music volume. The x-axis indicates the ratio of sensing signal volume and music volume. The default ratio is 1:1, and 0.6:1.4 is when music is roughly double in power, equivalently 7.3db louder. In summary, CoPlay is stable when training with agnostic music even in unseen volume. This could be attributed to the normalization in the definition of our amplitude-loss and recovery-loss.
Ablation of model architecture: In terms of the model design, we justify our choice by comparing the model performance under different model backbones. We considered three models namely CNN, RNN, and Wavenet with the same number of recursive layers. Fig. 5(a) shows that Wavenet slightly outperforms CNN and both surpass RNN. But CNN has a limited reception field so we chose Wavenet which is better at overcoming this issue with dilated convolution layers. Furthermore, the sample RNN uses LSTM layers which is less efficient to train in comparison.
Ablation of multi-task loss terms: The overall loss consists of sub-loss terms, each of which represents different objectives of optimization (section 3.3). By default, they are assigned equal weights. So, we ablate the weights to assess the performance when we emphasize on certain objective. In detail, recovery-loss is set to one by default, the same as other sub-losses. Hence, we also run the same ablation setup but increase the weight of recovery-loss from 1 to 1.2, 1.4, and 1.6, making it the largest sub-loss, i.e. making minimizing frequency distortion the most significant objective; this is a practical use case when certain sensing systems, like near-field tasks, can compromise the power for minimal frequency distortion. In Fig. 5(c), the error roughly goes down with the increase in weight, which serves as a reference for tuning this ratio for the downstream tasks in practice.
Ablation of sinc filter: Finally, we conduct an ablation to evaluate the effectiveness of the sinc filter as defined in section 3.3. It’s also the same experiment set up as the above ablation study. As shown in Fig. 5(d), both the sine and chirp models exhibit improved performance with the presence of sinc kernel, especially the sine wave model. We attribute this enhancement to the narrow target main lobe in the spectrum of sine, thus more side lobes are to be suppressed. The sinc kernel helps significantly reduce these lobes, resulting in a lower numerical loss. In summary, this ablation study validates the efficacy of the sinc filter, and we incorporate this essential layer in all of our experiments.
4.3. Spectral analysis
Training procedure in 3D frequency domain: Although the numerical loss can indicate the model performance, we seek a more comprehensive perspective by visualizing the model output in frequency spectrum. This helps confirm that the output aligns well with expectations. In Fig. 7, we take the FFT of a sample window output $\hat{x}$ and screenshot it with the increasing training epochs. The rise of the yellow peaks indicates that it converges to an 18kHz sine wave or 18k-20kHz FMCW as the training epoch increases. The side lobes are much weaker than the main lobes, which manifest good SNR. (Moreover, we further substantiate the SNR by qualitative study in section 5)
@@ 4.3. Spectral analysis
Figure 7. With the increasing of training epochs, the generated signal $\hat{x}$ grows to the expected shape, i.e. the target frequency rises while the unwanted noises are suppressed.
@@
Ranging profile with cross-correlation: Likewise, certain sensing algorithms employ cross-correlation (Xcorr) to obtain the ranging profile instead of utilizing FFT. Nevertheless, they basically use equivalent computation components: convolution or even DFT. For instance, an accelerated computation of Xcorr is to perform a frequency-domain Xcorr, which is faster than that in the time domain (Dr. Erol Kalkan, 2023), because shifting a signal in the time domain is equivalently scaling it in the frequency domain. Therefore, in addition to plotting the FFT spectrums of the generated signal, Fig. 8 shows the Xcorr correlation output for the cognitively-scaled signal. By rolling the signal in the time domain, the distinct peak in Xcorr output changes with the number of shifts as expected. In summary, this justifies that our adapted signal works well even if the receiver-side sensing algorithm uses Xcorr.
@@
Figure 8. Our generated has good cross-correlation property: the distinct peak moves accordingly when changing the number of shifts.
4.4. Computation efficiency
@@
We refer to CoPlay as a mixer algorithm, by expecting that our cognitive scaling can be applied to the buffered audio stream in the speaker. So, we run training and inference on small windows of audio clips. In this paper, we test with window size of 512, which is a common setting for acoustic sensing and also meets the computation efficiency requirement. To elaborate on the computation cost, our model has a size of 5.5 KB, forward/backward pass size of 0.98 MB, and an input size of 0.07 MB. The estimated total size of training is 1.05 MB only, which indicates the inference could easily fit into the RAM of commercial smartphones or other mobile devices. Despite that it was trained on an RTX 8000 GPU, it can run inference for roughly 625 fps on mobile devices, equivalently $1/625=1.6ms$ per frame. On the other hand, the buffer interval is roughly $512/48000Hz=10ms$, which leaves quite enough time for inference on one frame. Therefore, it can run in real time with no perceptible delay. Furthermore, the entire audio is actually pre-known by the speaker in many cases. For instance, Java sound API can load the audio file into memory and then buffer the sliding windows into input stream. That’s saying, the speaker can run the inference upfront and before playing the processed signal. Moreover, there is a lot of literature on model compression and inference acceleration to further improve the efficiency, which is beyond the scope of this work.
We also plot the roofline in Fig. 9, illustrating our model on RTX 3080 GPU in comparison with other popular models; ours is much more memory-efficient than even the mobile models like MobileNet, SqueezeNet, and DenseNet, because of its small number of parameters of dilated layers. We also show dashed rooflines from other popular mobile SoCs(System-on-a-Chip) and GPUs as a comparison of their peak performance and max memory bandwidth, which validates that our model would work on most chips, especially mobile chips like Snapdragon 888 and Kirin 9000.
Figure 9. Roofline plot
5. field study: respiration rate detection and hand gesture recognition
To better evaluate our method in practice beyond the micro-benchmark, we run our algorithm in two downstream acoustic sensing tasks: respiration rate detection and hand gesture recognition. These field studies demonstrate the power of CoPlay for improving the performance of various downstream tasks when there is interference from concurrent audio applications. The field studies are chosen to be controlled, to ablate-study the effects of our method, avoiding uncontrolled results from other factors like occlusion or range.
In summary, we run these 2 tasks with 2 types of sensing signals and 3 types of audio (piano, hot songs, and podcast) on 2 types of devices with 12 users in comparison with 4 baselines. In this section, we detail the tasks, evaluation metrics, hardware/software setups, and study design. Then we compare with 4 baselines of clipping, down-scaling, and no-music. The quantitative and qualitative results show that our method matches the sensing performance that has no music interference, while we have no issues as clipping and down-scaling have.
5.1. Tasks and evaluation metrics
In the field study, we evaluate our cognitive scaling mixer algorithm in two tasks: respiration rate detection and hand gesture recognition. In each task, we play inaudible ultrasound (sine or FMCW) from a mobile phone (IOS or Android) and then analyze received signals for recognition.
Respiration monitoring is a well-explored application in acoustic sensing, manifesting its high resolution that can detect subtle millimeter-level motions. The target is to get breath per minute BPM by detecting peaks in the waveform of chest movement during inhalations. Therefore, in the following results, we show the mean absolute error MAE of BPM and visualize chest movement in a waveform.
Hand gesture recognition is another important use case showing that acoustic sensing can leverage patterns in received signals for profound motion recognition. In this study, we classify three gestures: swipe, push, and pinch, which are essential in commercial gesture control systems. In the following result section, we show the classification accuracy and the confusion matrix as well.
5.2. Experiment setup
Hardware: We run acoustic sensing on two mobile phones: a Samsung smartphone and an iPhone, representing the mobile operation systems of Android and IOS, because their built-in algorithm handling overflow in the speaker mixer is different. As discussed above, Android turns to clip, IOS turns to scale down. Both phones are placed on the desk around 40cm away from the seated user facing the phone; one phone at a time.
@@ 5.2. Experiment setup
The ground truth for respiration detection is from the Go Direct Respiration Belt, a belt with a force sensor (Belt, 2023). The user wears it around the chest or abdomen. It also has a built-in breath rate calculation every 10 seconds. The sample window for the calculation is 30 seconds. To get the ground truth label for gesture recognition, we play a guiding video on a monitor and ask the user to follow the video to conduct specific gestures at a guided time every 3 seconds.
Data collection software: We developed our own data collection App for both Android and IOS. The recorder operates at a sample rate of 48k HZ. The recording format is 16-bit PCM. Both the transmitter and receiver are set to use the mono channel. And the volume of concurrent music is around 50db-60db. These are commonly supported audio parameters in mobile phones.
There is also a data collection program running on a laptop for synchronous ground truth data recording using the respiration monitoring belt. It calls the Python SDK of GoDirect (Belt, 2023) and runs as a Flask Restful API on a Macbook. The API is invoked by the data collection mobile App to start the recording of sensing and ground truth synchronously. The hand gesture guiding video also runs an API on the laptop to start playing the video synchronously.
Sensing algorithm: As for the receiver side sensing algorithm, we select two classic ones for the sine and FMCW signal separately. Note that we aim to show the effectiveness of CoPlay from the perspective of downstream task accuracy here, while many other factors could affect the accuracy, including the receiver-side sensing algorithm. So, the two selective ones are a baseline to show the relative effectiveness of our CoPlay V.S. baselines; Improving the receiver-side sensing algorithm is not the focus of this study.
As discussed above, almost all receiver-side algorithms are based on the Fourier transform in some way. Therefore, for the respiration rate detection, we play a sine wave signal at 18kHz and extract that frequency bin in the received signal, whose amplitude is proportional to the chest-to-device distance. Next, we apply peak detection on the amplitude waveform to get BPM. We use the basic Scipy peak detection function with default parameters; it finds all local maxima by simple comparison of neighboring values. Besides, to eliminate noisy motion other than the breath -for example, body or head rotation- we apply a band-pass filter before peak detection. The filter selects out 8-22Hz because the human breath rate is typically 12-18Hz. Specifically, we use the same window size of 30s and interval of 10s as the Vernier belt ground truth. The beginning ¡30s are padded.
@@
For the gesture recognition task, we transmit FMCW of 18k-20kHz and then dechirp the received signal with cross-correlation to get a range profile. At last, each window subtracts its previous window to eliminate the static environment and let the motions stand out. Note that nearby motions other than the hand would degrade sensing performance. So, we ask the users to try their best to stay still, since robustness under noise is not the focus of this study; we control these conditions to isolate the effectiveness of CoPlay only. Finally, the range profile is input into a simple CNN classifier (Document, 2023) as a single-channel 2D feature map. The CNN is a simple 2-conv-layer network with max pooling, ReLU, plus three linear layers at the end. Then by training with Adam optimizer and cross-entropy loss, we got a 3-class classifier. Besides, to avoid overfitting on our small dataset, we augment the feature map by adding Gaussian noise in the training dataloader. This regulation helps the training loss converge fast and stable. The classifier is trained in a leave-one-user-out manner and the following results are the average accuracy.
5.3. Study procedure
In the study, we first brief the participants on our IRB and study procedure without telling them the order and content of the audio to be played. Then we conduct 2 sessions of data collection; each has 6 sub-sessions. Upon finishing each sub-session, we ask the user to fill in a qualitative questionnaire about the music quality they heard. In total, each user collects 12 minutes of data and we have around 2 hours total data from 12 users.
For each user, the 2 sessions of data collection cover the two tasks of respiration and gesture detection separately. Within each session, there are 6 sub-sessions that are 6 minutes in total. The first three sessions run on the Android phone, playing 1) sensing signal only, no-music 2) audio + cognitively-scaled sensing 3) audio + sensing (cause clipping). The other three sessions run on the iPhone, playing the same audio except that the third case will cause down-scaling instead. Sub-session 1 and 3 are baselines to compare with our method in sub-session 2. Sub-sessions 2 and 3 use the same audio clip, which consists of samples from the datasets we used in the micro-benchmark section. It concatenates three 20-second clips from each of the three types of audio, including piano, hot songs, and speech. At the beginning of the first session, the user starts by putting on the belt and sitting in front of the device. Then they click the button on the phone screen to start recording. In between the sub-sessions, we switch the phone. Next, for the second session, they can take off the belt. Likewise, they will also click the button to start recording and playing the guiding video. Upon finishing every 3 sub-sessions, the user fills in the qualitative questionnaire about music quality and volume.
5.4. Study results with baselines
5.4.1. CoPlay V.S. Baselines (clipping and down-scaling)
@@ 5.4. Study results with baselines
First, we compare CoPlay with the baselines. They are no-music baseline, clip baseline, and two downscale baselines, measuring both the error of BPM and the accuracy of gesture recognition. As depicted in Fig. 10, our cognitively-scaled sensing with concurrent music achieves similar results as the sensing-with-no-music baseline; while the other three baselines exhibit inferior performance compared to the no-music. These results are an average of experiments across all devices, users, and audio types. Specifically, in the two downscale experiments, we reduce the power of sensing by half and leave the rest half for music. Additionally, the down_x2 refers to further scaling down the sensing signal to 1/4 of the original, resulting in a decrease of 12db in volume (sound pressure level is $20\log_{10}{power}$), then the music takes 3/4. Note that it is equivalent to double the sensing distance, as signal intensity is inversely proportional to the square of the distance from the source. In summary, both clipping and downscaling underperform our algorithm. Clipping emerges as the worst, even than large scaling down; while downscaling would get worse quadratically with increasing distance in practice, although the drop is less noticeable when compared with the worst case from clipping in our experiment. In the subsequent section, we will elucidate these observations by visually analyzing their breakdowns.
@@
Figure 10. The performance of our CoPlay V.S. 4 baselines, for task (a) and task (b) separately.
5.4.2. Visualization and error analysis
Figure 11. The amplitude of detected breath rhythm of the three scenarios: no-music, CoPlay, and clip(Android) or downscale(IOS).
Fig. 11 depicts the amplitude of detected breath rhythm of the three scenarios: no-music, CoPlay, and clip(Android) or downscale (IOS). To recap, the amplitude of breath rhythm is the normalized amplitude of FFT at 18k, named $f_{18k}$, which is roughly proportional to the chest-to-device distance. First, it is obvious that the no-music scenario for both Android and IOS is stronger than our cognitively scaled signal. While clipping also keeps the received amplitude strong, it is significantly noisy because it is distorted by the mixer at the transmitter. Relatively, CoPlay and no-music scenarios are clear periodic waveforms. Secondly, the downscale case is almost two times weaker than the no-music case, while CoPlay is only 15% weaker. Besides, its crests and troughs are not as observable as the others, which is ineffective for peak detection. The ground truth BPM of the two selected samples are a) 18.5 and b) 18.9. As shown in the legend in the figures, CoPlay and the no-music baseline are close to the ground truth BPM while clip and downscale have a larger error. The dots along each line are the individual breath selected by the peak detection algorithm.
In Fig. 12, to take a closer look at the visualization of our output, we zoom in on the first line plot of CoPlay and plot it with the normalized ground truth in the background. The dashed line is the force sensor recording, from which the Vernier Belt derives the BPM. The peaks align well even though the overall amplitude rises and falls, which substantiates the effectiveness of CoPlay.
Figure 12. A closer look at the detected breath rhythm by our method, in comparison with the ground truth by the tension-force-sensor belt.
Furthermore, we take a closer look at the errors of gesture classification by calculating the confusion matrix that plots accuracy for each class. Since our method and no-music baseline are 100% accurate, Fig. 13 lists the confusion matrix of the downscale baseline and the clip baseline. As expected, wireless sensing is more sensitive to vertical motion, i.e. pushing close to or away from the device, because the distance of reflection traveling in the air changes dramatically. So, the errors are high in the swipe and pinch rows/columns in the confusion matrix, while push is almost always recognized accurately.
@@ 5.4.2. Visualization and error analysis
Figure 13. Confusion matrix of gesture classification.
5.4.3. Qualitative study
Table. 2 is the result of the qualitative study. They are the playback quality of each sub-session per level of buzzing noise, loudness, and delays. The scores are the average of all users’ answers, asking for a 1 - 5 scale (integer) per question:
  • Q.1. Did you hear a buzzing noise?
  • Q.2. Did you hear the music at adequate volume?
  • Q.3. Did you perceive delay or discontinuity?
The score of Q.1 justifies that clipping does cause audible noise, basically buzzing, while CoPlay and down-scaling do not. Q.2 manifests that down-scaling hurts the power of audio signals, resulting in the users being unable to hear the audio at a sufficient volume. To elaborate, in the study, we calibrated the system volume of the phone to 50db, a common loudness of music playback or a nearby conversational talk. The volume was measured by another smartphone App placed next to the study device. When down-scaling happens, the music volume reduces by 6db per each half down-scaling. Q.3 is to prove that the computation overhead of CoPlay does not cause any discontinuity or delay between the chunks of processed clips.
@@ 5.4.3. Qualitative study
Q.1 (5 = buzzing) 4.72 1.13 1.22
Q.2 (5 = loud) 4.14 1.12 4.49
Q.3 (5 = delay) - - 1
Table 2. Average qualitative study results (1-5 scale).
6. discussion and future work
In this work, our system handles two types of sensing signals on two downstream applications. In the future, in order to explore the full extent of applicable scenarios and thoroughly assess the versatility of the proposed approach, it would be prudent to engage in a more expansive array of field studies and sensing signals. These studies could extend beyond the initial scope of simple tasks and delve into more intricate downstream tasks such as more fine-grained gesture recognition. Specifically, evaluating the algorithm’s performance in challenging environments, characterized by varying degrees of interference (like motion, external noise, etc), could yield valuable insights into its robustness and effectiveness under diverse conditions.
Moreover, we recognize a potential escalation in computational costs, particularly when dealing with multiple channels. Presently, our assumptions are based on the notion that devices with multiple channels duplicate the same signal into each channel. However, in the realm of 3D spatial audio applications, there exists the possibility of having slightly different signals per channel. Although our algorithm remains non-intrusive to the music itself, the need for custom cognitive scaling for each channel introduces an increase in computational demands. It would be worthy to delve deep into evaluating the computation cost of our algorithm in these more realistic scenarios and apply state-of-the-art acceleration methods from the perspectives of both the model compression and on-device streaming data processing.
Furthermore, we aim to delve deeper into investigating the behavior of speaker mixers through the execution of real-world experiments across a more diverse range of devices. This expanded investigation seeks to provide a more comprehensive understanding of how speaker mixers function in various technological contexts. Besides, the variant of the number of sensors and their positions are also relevant factors remained to be investigated. For example, smart speakers like Amazon Echo Dot and Google Home have much more powerful speakers and the full stack runs on unique IoT operating systems, the combination of which could bring distinct insights to the scope of our concurrent music problem.
@@
Amplitude modulation (AM) is a modulation technique where the amplitude of a carrier signal is varied in proportion to the instantaneous amplitude of a modulating signal. Here is a simple case and the formulas of how the side frequencies derivative(Wikipedia, 2023).
@@
(9) $A_{m}\cos(2\pi f_{c}t)\cos(2\pi f_{m}t)$
(10) $=$ $\frac{A_{m}}{2}[\cos(2\pi(f_{c}+f_{m})t)+\cos(2\pi(f_{c}-f_{m})t)]$
(11) $=$ $\frac{A_{c}}{2}\frac{A_{m}}{A_{c}}[\cos(2\pi(f_{c}+f_{m})t)+\cos(2\pi(f_{c}-f_{m})t)]$

### v3

@@ IV-C Spectral analysis
Figure 5: Ablation study results.
@@
Figure 6: Generated signal $\hat{x}$ grows as expected as training epochs increase: target frequency rises, noises are suppressed.
Figure 7: Roofline plot.
Figure 8: The amplitude of detected breath rhythm of the three scenarios: no-music, CoPlay, and clip(Android) or downscale(IOS).
@@
Ranging profile with cross-correlation: Likewise, certain sensing algorithms employ cross-correlation (Xcorr) to obtain the ranging profile instead of utilizing FFT. Nevertheless, they basically use equivalent computation components: convolution or even DFT. For instance, an accelerated computation of Xcorr is to perform a frequency-domain Xcorr, which is faster than that in the time domain [38], because shifting a signal in the time domain is equivalently scaling it in the frequency domain. Therefore, in addition to plotting the FFT spectrums of the generated signal, Fig. 9 shows the Xcorr correlation output for the cognitively-scaled signal. By rolling the signal in the time domain, the distinct peak in Xcorr output changes with the number of shifts as expected. In summary, this justifies that our adapted signal works well even if the receiver-side sensing algorithm uses Xcorr.
@@
Figure 9: Generated signal has good cross-correlation property: the distinct peak moves with change of shifts.
IV-D Computation efficiency
@@
CoPlay should function as a mixer algorithm for buffered audio stream in the speaker, processing small windows of audio clips. In this paper, window size is 512, common for acoustic sensing and efficiency. Our model size is as compact as 5.5 KB, with forward/backward pass size of 0.98 MB, and an input size of 0.07 MB. The estimated total size of training is 1.05 MB, easily fitting into the RAM of commercial smartphones or other mobile devices during inference. Despite training on an RTX 8000 GPU, inference runs for roughly 625 fps on mobile devices, equivalently $1/625=1.6ms$ per frame. So, a buffer interval around $512/48000Hz=10ms$ could leave quite enough time for inference on one frame, making it real-time with no perceptible delay. Furthermore, the entire audio is pre-known by the speaker in many cases, so it can process the signal upfront before playing. Fig. 8 also shows our superior efficiency compared to other popular models on RTX 3080; we even outperform mobile models like MobileNet, SqueezeNet, and DenseNet, due to our small dilated layers. It’s equivalently 10X more power-efficient than background Apps. The dashed rooflines from other popular mobile SoCs(System-on-a-Chip) and GPUs compare the peak performance and memory bandwidth, validating our viability on most mobile chips, including Snapdragon 888 and Kirin 9000. Moreover, literature on model compression and acceleration could further improve the efficiency, which is beyond the scope of this work.
V field study
To validate CoPlay’s efficacy in real-world beyond the micro-benchmark, we run two downstream acoustic sensing tasks: respiration rate detection and hand gesture recognition, with concurrent audio. The study is across 2 tasks, 2 sensing signals, 3 audio types, and 2 device types with 12 users, compared against 3 baselines. Experiments are controlled to isolate CoPlay’s effects, avoiding confounding factors like occlusion or range. Moreover, since CoPlay process signals before playback in speaker, downstream factors (like environmental interference, audible leakage and nonlinearity) wouldn’t impact CoPlay upstream and can be addressed via other existing methods such as filtering [8, 6].
V-A Field study tasks and evaluation metrics
1) Respiration monitoring is a well-explored case, manifesting the ability of millimeter-level detection of acoustic sensing. It measures breath per minute (BPM) by detecting waveform peaks of chest movement during inhalations. So, our results report mean absolute error (MAE) of BPM alongside waveform visualizations. 2) Hand gesture recognition highlights motion recognition capabilities of acoustic sensing. Here we classify three gestures: swipe, push, and pinch, which are essential in gesture control systems. Our results report classification accuracy alongside the confusion matrix.
V-B Field study experiment setup
@@
Hardware: We test with two mobile phones: a Samsung S10 phone and an iPhone 13, representing Android and IOS, whose current solutions to mixer overflow are different: Android turns to clip, IOS turns to scale down. Phones are placed on the desk around 40cm away from the seated user facing front; one at a time. For respiration detection, ground truth is from the Go Direct Belt with a force sensor [39]. It has a built-in breath rate calculation every 10 seconds with a window size of 30 seconds. For gesture recognition, users follow a guided video displaying gestures every 3 seconds.
Software: We developed a custom data collection app for both Android and iOS. The recorder operates at 48kHz with 16-bit PCM, using a mono channel for both transmission and reception. Concurrent music plays at 50-60dB, aligning with typical mobile audio settings. A laptop runs a data collection program for synchronized ground truth recording from the belt SDK [39] or gesture guiding video playback. It operates as a Flask RESTful API on a MacBook, triggered by the mobile app for synchronous sensing and ground truth logging.
@@
Sensing algorithm: We use two classic receiver-side algorithms for sine and FMCW separately. While various factors affect accuracy, including the sensing algorithm itself, our focus is on demonstrating CoPlay’s effectiveness relative to baselines, not optimizing the receiver-side algorithm. 1) For respiration rate detection, we use an 18kHz sine wave, extracting its frequency bin from the received signal, whose amplitude corresponds to chest-to-device distance; then we apply Scipy peak detection function with default parameters to estimate BPM. Besides, a band-pass filter (8-22Hz) before peak detection can filter out non-breathing motion like body rotation, since typical human breath rate is 12-18Hz. The analysis follows Vernier belt’s 30s window and 10s interval, with padding in initial 30s data. 2) For gesture recognition, we transmit FMCW of 18k-20kHz and then dechirp the received signal with cross-correlation to get a range profile. Each window subtracts its previous one to eliminate the static environment and let the motions stand out. To minimize interference from unintended nearby movements, users are instructed to stay still, isolating CoPlay’s effectiveness since robustness under noise is not the focus of this study. Finally, the range profile, as a single-channel 2D feature map, is input into a simple CNN classifier [40], which is a 2-conv-layer network with max pooling, ReLU, plus three linear layers. Then by training with Adam optimizer and cross-entropy loss, we got a 3-class classifier. Besides, to avoid overfitting on our small dataset, we augment the feature map by adding Gaussian noise. Training follows a leave-one-user-out scheme, and reported results are the average accuracy.
V-C Study procedure
Participants are first briefed on the IRB and study procedure without disclosing the audio order or content to play. Each user had 2 sessions (one for respiration, one for gestures), with 6 sub-sessions per session, totaling 12 minutes, i.e. over 2 hours for all 12 users. Each session included three sub-sessions on Android phone: 1) sensing signal only, no-music, 2) audio + cognitively-scaled sensing, and 3) audio + sensing (causing clipping). Another three sub-sessions ran on iPhone similarly except that the third case will cause down-scaling instead. Sub-sessions 1 and 3 served as baselines to compare with CoPlay in sub-session 2. Audio clips were from datasets used in the micro-benchmark, featuring three 20-second samples of piano, hot songs, and speech. Users started recording, switched phones and belt between sub-sessions, and then filled out qualitative questionnaires after every three sub-sessions.
V-D Field study results with baselines
@@
CoPlay V.S. Baselines (clipping and down-scaling). The four baselines are no-music baseline, clipping baseline, and two downscale baselines, measuring the MAE of BPM and gesture recognition accuracy. As depicted in Fig. 10, CoPlay with concurrent music achieves similar results as the no-music baseline; while all the other three baselines exhibit inferior performance than the no-music. These findings are an average of across devices, users, and audio types. In down-scaling, sensing power is reduced by half with the remaining half allocated to music. The down_x2 further scales to 1/4 of the original, lowering volume by 12db (as sound pressure level is $20\log_{10}{power}$), then the music takes 3/4; it is also equivalent to double the sensing distance, as signal intensity is inversely proportional to the square of the distance from the source. Overall, both clipping and downscaling underperform compared to CoPlay. Clipping is the worst, even more than aggressive downscaling, while downscaling may degrade quadratically with increasing distance. The next section will elucidate these observations by a visual breakdown.
@@
Figure 10: Compared with sensing with no music, the baselines degrade sensing performance, while CoPlay does not.
Visualization and error analysis. Fig. 8 depicts the amplitude of detected breath rhythm of the three scenarios: no-music, CoPlay, and clip(Android) or downscale (IOS). To recap, the amplitude of breath rhythm is the normalized FFT amplitude at 18kHz, named $f_{18k}$, roughly proportional to the chest-to-device distance. First, it is obvious that the no-music scenario for both Android and IOS is the strongest. While clipping maintains a high amplitude, it is significantly noisy due to distortion. In contrast, CoPlay and no-music produce clear periodic waveforms. Secondly, downscaling weakens the signal by nearly 50%, while CoPlay reduces it by only 15%; Its crests and troughs are also less distinct, making peak detection ineffective. The ground truth BPM of the two selected samples are 18.9 and 18.5. As shown in figure legend, CoPlay and the no-music baseline are close to the ground truth BPM while the others have a larger error. The dots along each line are the individual breath selected by the peak detection algorithm. Fig. 11 further zooms in on the first CoPlay plot, overlaying it with the ground-truth dashed line (the normalized raw data from the force sensor). The peaks align well despite amplitude variations, validating CoPlay’s effectiveness.
Figure 11: The detected breath of CoPlay V.S. ground truth.
Similarly, we take a closer look at the errors of gesture classification using confusion matrix. Since CoPlay and the no-music baseline achieve 100% accuracy, Fig. 12 only presents results of downscaling and clipping. As expected, wireless sensing is more sensitive to vertical motions like pushing, where reflection distance changes significantly. Thus, errors are higher for swipe and pinch, while push is almost always accurately recognized.
Qualitative study. Table. II is the qualitative study result, evaluating the playback quality of each sub-session per level of buzzing noise, loudness, and delays. The scores are the average of all users, in a 1-5 scale (integer) per question: Q1) Did you hear a buzzing noise? Q2) Did you hear the music at adequate volume? Q3) Did you perceive delay or discontinuity? In general, the results justify that CoPlay leaves the music untouched. While Q1 manifests that clipping causes audible noise, basically buzzing. Q2 shows that downscaling hurts audio power, making music less audible to users. To elaborate, in the study, we calibrated the system volume of the phone to 50db, a common loudness of music playback or a nearby conversational talk. The volume was measured by another smartphone App placed next to the study device. When downscaling happens, the music volume reduces by 6db per each half downscaling. Q3 is to prove that the computation overhead of CoPlay does not cause any discontinuity or delay between the chunks of processed clips.
@@
Figure 12: Confusion matrix of gesture classification.
@@
Q.1 ($\downarrow$5 = buzzing) 4.72 1.13 1.22
Q.2 ($\uparrow$5 = loud) 4.14 1.12 4.49
Q.3 ($\downarrow$5 = delay) - - 1
TABLE II: Qualitative study results (1-5 scale): CoPlay do not degrade music but clipping and downscaling do.
VI discussion and future work
This work explores 2 sensing signals in 2 applications. Future studies should explore a more expansive array of cases to thoroughly assess the versatility of the proposed approach. It should engage more intricate downstream tasks, like fine-grained gesture recognition, on diverse hardware.
We acknowledge a potential increase in computational costs, particularly for multiple channels. Although we assume multiple channels carry duplicate signals, 3D spatial audio may leverage slight variations requiring custom scaling and add to the computation load. Future work should evaluate this cost in realistic scenarios and explore acceleration methods like model compression and acceleration, also in comparison with other sensors’ resource cost, like an MCU.
We aim to further explore speaker mixer behavior through real-world experiments on diverse devices, since factors like sensor count and placement might warrant investigation. Smart speakers like Amazon Echo Dot and Google Home, with their custom hardware and OS, may offer distinct insights into the concurrent music problem.
@@
Amplitude modulation (AM) is a modulation technique where the amplitude of a carrier signal is varied in proportion to the instantaneous amplitude of a modulating signal. Here is a simple case and the formulas of how the side frequencies derivative[41].
@@
$A_{m}\cos(2\pi f_{c}t)\cos(2\pi f_{m}t)$ (9)
$=$ $\frac{A_{m}}{2}[\cos(2\pi(f_{c}+f_{m})t)+\cos(2\pi(f_{c}-f_{m})t)]$ (10)
$=$ $\frac{A_{c}}{2}\frac{A_{m}}{A_{c}}[\cos(2\pi(f_{c}+f_{m})t)+\cos(2\pi(f_{c}-f_{m})t)]$ (11)